OpenStudy (anonymous):

A cash register contains only five dollar and twenty dollar bills. It contains three times as many five dollar bills as twenty dollar bills and the total amount of money in the cash register is $735. How many five dollar bills are in the cash register?

OpenStudy (anonymous):

Let x=number of $5, y= number of $20 5x+20y=735

OpenStudy (anonymous):

Let x be how many $20s you have in the register Then you can set up the equation as such \[3x($5) + x($20) = 735\] Solve for x and that will give you how many $20s you have in your reg. Then using that, multiply "x" by 3 to find out how many $5s you have (since you have 3 times as many five dollar bills as twenty)

OpenStudy (anonymous):

Let x=number of $5, y= number of $20 5x+20y=735 x=3y 5(3y)+20y=735 Use substitution to solve.

OpenStudy (anonymous):

when you do that, you get x=63 and y=21

OpenStudy (anonymous):

so there are 63 five dollar bills